OpenStudy (anonymous):

Solve three times square root of x minus 3 end root minus 2 equals 16 (3)sqrt(x-3) -2 = 16

OpenStudy (anonymous):

Ok, so we have to solve for x. Add +2 on the left and right side, and then divide each side by three. Ok so far?

OpenStudy (anonymous):

Then we will have \[\sqrt{x-3} = \frac{ 16+2 }{ 3 }\], or \[\sqrt{x-3} = 6\]

OpenStudy (anonymous):

Thank you very much!!! :))

OpenStudy (anonymous):

In order to solve for x, we have to get rid of the square root. Hence, we do the opposite operation, namely squaring both sides, which implies \[x-3 = 6^2\]. Ok there you go! Glad I could help :)